# Learning phase transitions by siamese neural network

### v1

@@ Abstract
The wide application of machine learning (ML) techniques in statistics physics has presented new avenues for research in this field. In this paper, we introduce a semi-supervised learning method based on Siamese Neural Networks (SNN), trying to explore the potential of neural network (NN) in the study of critical behaviors beyond the approaches of supervised and unsupervised learning. By focusing on the (1+1) dimensional bond directed percolation (DP) model of nonequilibrium phase transition, we use the SNN to predict the critical values and critical exponents of the system. Different from traditional ML methods, the input of SNN is a set of configuration data pairs and the output prediction is similarity, which prompts to find an anchor point of data for pair comparison during the test. In our study, during test we set different bond probability $p$ as anchors, and discuss the impact of the configurations at this anchors on predictions. More, we use an iterative method to find the optimal training interval to make the algorithm more efficient, and the prediction results are comparable to other ML methods.
Introduction
@@
The mainly types of ML method applied to study phase transitions are supervised learning [5, 8, 9, 10, 11], unsupervised learning [6, 12, 13, 14, 15, 16, 11], and semi-supervised learning  [17, 18, 19, 20, 21]. For supervised learning, the input data for training needs to be labeled, which commonly used to identify or classify the phases of matter. For unsupervised learning, the labeling of the order parameter is not required, such as the methods of principal component analysis (PCA)[22, 12, 16], t-distributed stochastic neighbor embedding (T-SNE) [23] or nonlinear autoencoder (AE) [23, 11, 12, 16], for clustering and dimensionality reduction. Recently, semi-supervised learning, especially transfer learning (TL) also has been proved as a good choice to detect critical points and calculate critical exponents [20], even determining the type of phase transition [21]. With the input of mixing both labeled and unlabeled data, methods of TL can transform unlabeled data in target domain into labeled data in source domain by training, which only use a limited labeled set for the purposes of prediction. Although the methods of ML are dazzling, all of them are not fully perfect. For example, supervised learning must time-consumingly label the data; unsupervised learning is hard to fully identify the dynamic features with limited data; semi-supervised learning is very sensitive to the input, which always need to be preprocessed. Therefore, it is still imperative to find a more complete and efficient method of ML to study phase transitions.
@@
Siamese neural network (SNN) [24, 25, 26], also called twin NN, is a specialized architecture usually consisting of two parallel identical neural networks (NN) that share their weight for the purposes of assessing or comparing the similarity between two input objects. SNN is not simply classified into supervised learning, unsupervised learning or semi-supervised learning, but a flexible structure that fill them in. It is widely used in computer vision for face recognition [27], target tracking [28] and image retrieval [29], and also shows its power in natural language processing (NLP) for text matching tasks [30]. In the study of phase transition, Patel $etal$ [31] introduced a SNN technique of unsupervised learning to identify phase boundaries in Ising-type systems and Rydberg atomic arrays, showing its ability to learn about multiple phases without knowing about their existence. As the preliminary attempt, SNN is proved as a valuable tool for discovering new and unfamiliar phases of matter, and worth for further application and exploration.
@@
Inspired by this, we apply a new SNN algorithm based on semi-supervised learning to study the critical behaviors of phase transition. In this method, we divide input configurations to two separate sets for training. By choosing an anchor for configurations, the SNN would determine whether this samples belong to the same set through training, which gives the power to judge phases and find critical point. For test, we focus on the model of (1+1)-dimensional bond directed percolation (DP) [32, 33, 34], belonging to the most important universality class of nonequilibrium phase transitions, the DP universality class [35, 36, 37]. We not only predict the critical point of the DP model through the SNN, but also calculate the critical exponent by the data collapse method [38, 39]. Further, we discuss the impact of anchor point selection on results, and find the optimal interval of training set to achieve more accurate predictions. By doing so, we hope to enhance the effectiveness and reliability of the SNN as a new tool in the study of phase transitions.
@@
The remaining content of this paper is structured in the following manner. In Section 1, we have introduced the directed percolation model. Section 2 gives the methods of SNN and the data set for study. Section 3 shows the estimate critical probabilies of SNN and also the prediction of spatial correlation exponent $\nu_{\perp}$. In Section 4, we discuss the case of chosen archor in critical region, and the impact of the training set on the results. Section 5 is a summary of this work.
1 Model
@@ 1.1 The directed percolation model
As the most prominent universality class of absorbing phase transition, DP describes percolation of one direction along that the system falls into inactive state where the time evolution stops. It is utilized to investigate various phenomena including the propagation of epidemics [40], forest fires [41], and traffic flow  [42], etc. In past decades, a various of methods has been applied to study DP models, such as mean-field theory [43], renormalization groups [44], field theory approach [45] and numerical simulations [46] and so on. Some ML approaches have also been tried recently and shown their own strengths [20, 47]. Since the models of DP universality class include unitary or binary random reaction processes, which may contain diffusion or non-diffusion motions, only the critical exponents are common and the position of the phase transition may vary. In our work, we use a SNN based algorithm to study the $(1+1)$-dimensional bond DP, one of the simplest cases of DP universality class.
@@
$\left\{\begin{array}[]{rcl}\mbox{self-destruction:}&A\longrightarrow\varnothing,\\ \mbox{diffusion:}&\varnothing+A\longrightarrow A+\varnothing,\\ \mbox{offspring production:}&A\longrightarrow 2A,\\ \mbox{coagulation:}&2A\longrightarrow A.\\ \end{array}\right.$ (1)
@@
Figure 1: Configurations of the (1+1) dimensional bond DP corresponding to three different bond probabilities $p=0.5,0.6447$ and $0.8$ respectively.
@@
where $s_{i}(t)$ is the state of site $i$ at time $t$ and $p$ is the bond probability of a connection between two adjacent sites. $z_{i}^{\pm}\in(0,1)$ is a random number taken from an uniform distribution. From the definitions above, the configuration of the system can be determined by the states of all sites.
The models of DP universality class also has their own independent critical exponents. The order parameter of the (1+1)-dimensional bond DP close to the critical point $p_{c}$ can be expressed by the steady-state particle density $\rho_{a}$ according to the critical exponent $\beta$ as [35],
@@
In this paper, we use the the full configurations of the (1+1)-dimensional DP shown in Fig. 1 as input to the SNN algorithm. To generate the configurations, we use periodic boundary conditions, and originally starts with fully occupied lattice (explained in our previous study [20]). The characteristic temporal length $t_{c}\sim L^{z/d}$ is set by the dynamical exponent $z=1.580(1)$ and spatial dimension $d=1$, to ensure the stable results [35].
@@ 2.1 The Siamese Neural Network (SNN) method
Figure 2: The structure of SNN
@@
The specific framework of SNN used in our work to study phase transition of the (1+1)-dimensional DP is shown in the Figure 2. We use the raw configurations of $L=8,16,32,48$ and $64$ as input, and the details of dataset selection will be given in Section 2.2. The input is a set of data pairs $(S_{i},S_{j})$, $S_{i}$ enters one branch in the framework, and $S_{j}$ enters the other. $S_{i}$ and $S_{j}$ are two different configuration samples of the same type of data. Taking the branch with $S_{i}$ as an example, the input $S_{i}$ is flattened into one-dimensional data and fed into an fully connected (FC) layer of $500$ neurons with a $ReLu$ activation function. Here we also set a Batch Normalization (BN) to avoid over-fitting and speed up the NN to optimize its parameters. The output of this FC layer is a feature representation $f_{1}$ of length $50$. Therefore, the entire calculation process of the first branch can be expressed as function $f_{1}(S_{i})$. Similarly, the second branch with $S_{j}$ has the same process, and the feature representation is recorded as $f_{2}(S_{j})$. Note that these two branches shares the weights for learning.
@@
The shared weights enable embedded data to be in the same space from any branches. Therefore, it allows to calculate the “distance" between two the feature representations $f_{1}(S_{i})$ and $f_{2}(S_{j})$, to compare the similarity of $S_{i}$ and $S_{j}$. Here we define the “distance" by $L_{1}$ norm. Then this information of “distance" is fed into a dense layer of FC with the numbers of neurons $50$ to $1$, and this process is also attached with a $ReLu$ activation function and BN. By applying a $Sigmoid$ activation function to the output of FC, the result is converted into a probability of $[0,1]$, which is regarded as classification score $y$, for similarity.
@@ 2.2 Data sets of the SNN
For input data of the SNN, we use the full configurations of (1+1) dimension DP (time evolution step as the characteristic temporal length.). In each lattice size $L=8,16,32,48$ and $64$, we select $101$ values of $p\in[0,1]$ uniformly and generate $1000$ samples of configuration at each $p$. To minimize human intervention during training, we label the samples of configuration far away from the critical regime, i.e., the input configurations $S_{i}$ and $S_{j}$ are generated in the range of $p\in[0,0.1]\cup[0.9,1]$. The idea of labeling here is very different from the input of TL in Ref. [20]. The labels “1" or “0" indicates the input data pair $S_{i}$ and $S_{j}$ belong to the same or different phases to ensure maximum similarity or dissimilarity between them. Specifically, if $S_{i}$ and $S_{j}$ both belong to $[0,0.1]$ (or $[0.9,1]$), the input $(S_{i},S_{j})$ is labeled as “1". Or $S_{i}$ and $S_{j}$ are in the range of $[0,0.1]$ and $[0.9,1]$ respectively, the data pair is labeled as “0". For training, the loss function is defined as:
@@
where $s$ is the similarity of ($S_{i}$, $S_{j}$) , and $y$ is the label, ranging from 0 to 1. The training epoch is set to $5000$, and our work in on python 3.10 and RTX 4090 Gpu partform with tensorflow 2.14.0 library.
For the process of test, a reference point, called "anchor" point, is selected to evaluate the similarity $s$ or dissimilarity $(1-s)$ of all configurations. The anchor $p_{a}$ can be selected arbitrarily within the range of $p\in[0,1]$. By taking the samples of configuration at the anchor $p_{a}$ as $S_{j}$ and the samples of the training set as $S_{i}$, in this way the similarity between the samples at the anchor and the training samples can be estimated through the SNN. And in detail the dependence of SNN results on anchor point will be discussed in Section 3.2. Then the output of SNN, $P_{0}$, can be expressed as two curves within bond probability $p\in[0,1]$: the curve $s$ of similarity and $1-s$ of dissimilarity. The intersection of these two curves or $p$ at $P_{0}=50\%$ represents the critical point of phase transition in the system.

### v2

@@ Abstract
The wide application of machine learning (ML) techniques in statistics physics has presented new avenues for research in this field. In this paper, we introduce a semi-supervised learning method based on Siamese Neural Networks (SNN), trying to explore the potential of neural network (NN) in the study of critical behaviors beyond the approaches of supervised and unsupervised learning. By focusing on the (1+1) dimensional bond directed percolation (DP) model of nonequilibrium phase transition and the 2 dimensional Ising model of equilibrium phase transition, we use the SNN to predict the critical values and critical exponents of the systems. Different from traditional ML methods, the input of SNN is a set of configuration data pairs and the output prediction is similarity, which prompts to find an anchor point of data for pair comparison during the test. In our study, during test we set different bond probability $p$ or temperature $T$ as anchors, and discuss the impact of the configurations at this anchors on predictions. In addition, we use an iterative method to find the optimal training interval to make the algorithm more efficient, and the prediction results are comparable to other ML methods.
keywords:
Machine learning , Siamese Neural Network , Directed percolation , Ising model , Phase transitions
journal: Physica A-Statistical Mechanics and Its Applications
1 Introduction
@@
The mainly types of ML method applied to study phase transitions are supervised learning [5, 8, 9, 10, 11, 12], unsupervised learning [6, 13, 14, 15, 16, 17, 11, 18], and semi-supervised learning  [19, 20, 21, 22, 23]. For supervised learning, the input data for training needs to be labeled, which commonly used to identify or classify the phases of matter. For unsupervised learning, the labeling of the order parameter is not required, such as the methods of principal component analysis (PCA)[6, 13, 17, 18], t-distributed stochastic neighbor embedding (T-SNE) [24] or nonlinear autoencoder (AE) [24, 11, 13, 17, 25], for clustering and dimensionality reduction. Recently, semi-supervised learning, especially transfer learning (TL) also has been proved as a good choice to detect critical points and calculate critical exponents [22], even determining the type of phase transition [23]. With the input of mixing both labeled and unlabeled data, methods of TL can transform unlabeled data in target domain into labeled data in source domain by training, which only use a limited labeled set for the purposes of prediction. Although the methods of ML are dazzling, all of them are not fully perfect. For example, supervised learning must time-consumingly label the data; unsupervised learning is hard to fully identify the dynamic features with limited data; semi-supervised learning is very sensitive to the input, which always need to be preprocessed. Therefore, it is still imperative to find a more complete and efficient method of ML to study phase transitions.
@@
Siamese neural network (SNN) [26, 27, 28], also called twin NN, is a specialized architecture usually consisting of two parallel identical neural networks (NN) that share their weight for the purposes of assessing or comparing the similarity between two input objects. SNN is not simply classified into supervised learning, unsupervised learning or semi-supervised learning, but a flexible structure that fill them in. It is widely used in computer vision for image recognition [29], target tracking [30] and image retrieval [31], and also shows its power in natural language processing (NLP) for text matching tasks [32]. In the study of phase transition, Patel etaletal [33] introduced a SNN technique of unsupervised learning to identify phase boundaries in Ising-type systems and Rydberg atomic arrays, showing its ability to learn about multiple phases without knowing about their existence. The SNN of convolutional neural network(CNN) architectures has also been used to investigate phase transitions based on polarising microscopic textures of liquid crystals, showing high accuracy for phase distinguishing [34]. As the preliminary attempt, SNN is proved as a valuable tool for discovering new and unfamiliar phases of matter, and worth for further application and exploration. Unlike general ML methods used on phase transition models, the SNN is not intended to predict which phase an input configuration belongs to, but judges whether different configurations belong to a same phase. Therefore, the output of the SNN is a measure of similarity, not a label associated with the input configuration. The SNN compares the configuration of unknown labels with the configuration of specific anchor points, transforming the process of assigning predicted labels into judging similarity. This means the selection of training and test intervals becomes very flexible, allowing SNNs to handle not only two-phase but also multi-phase complex statistical models [35], with a significant improvement in computing efficiency.
@@
Inspired by this, we apply a new SNN algorithm based on semi-supervised learning to study the critical behaviors of phase transition. In this method, we divide input configurations to two separate sets for training. By choosing an anchor for configurations, the SNN would determine whether this samples belong to the same set through training, which gives the power to judge phases and find critical point. For test, we focus on two key models. The first is the (1+1) dimensional bond-directed percolation (DP) [36, 37, 38], which belongs to the DP universality class, a prominent class of non-equilibrium phase transitions [39, 40, 41]. The second model is the 2 dimensional Ising model [42, 43], which belongs to the Ising universality class. We not only predict the critical point of the DP model through the SNN, but also calculate the critical exponent by the data collapse method [44, 45]. Further, we discuss the impact of anchor point selection on results, and find the optimal interval of training set to achieve more accurate predictions. By doing so, we hope to enhance the effectiveness and reliability of the SNN as a new tool in the study of phase transitions.
@@
The remaining content of this paper is structured in the following manner. In Section 2, we have introduced the two models. Section 3 gives the methods of SNN and the data set for study. Section 4 shows the estimate critical probabilities of SNN and also the prediction of spatial correlation exponent $\nu_{\perp}$. In Section 5, we discuss the case of chosen archor in critical region, and the impact of the training set on the results. Section 6 is a summary of this work.
2 Models
@@ 2.1 The directed percolation model
As the most prominent universality class of absorbing phase transition, DP describes percolation of one direction along that the system falls into inactive state where the time evolution stops. It is used to investigate various phenomena, including the propagation of epidemics [46], forest fires [47], and traffic flow  [48], etc. In past decades, various methods have been applied to study DP models, such as mean-field theory [49], renormalization groups [50], field theory approach [51] and numerical simulations [52] and so on. Some ML approaches have also been tried recently and shown their own strengths [22, 53]. Since the models of DP universality class include unitary or binary random reaction processes, which may contain diffusion or non-diffusion motions, only the critical exponents are common and the position of the phase transition may vary. In our work, we use a SNN based algorithm to study the $(1+1)$-dimensional bond DP, one of the simplest cases of DP universality class.
@@
[…]
@@
Figure 1: Configurations of the (1+1) dimensional bond DP corresponding to three different bond probabilities $p=0.5,0.6447$ and $0.8$ respectively. The Black sites represent occupied sites, while the white areas indicate empty regions. The system configuration corresponds to $L=500$ with $t=500$ time steps. The vertical and horizontal arrows denote the temporal evolution direction and spatial dimensions, respectively. The periodic boundary conditions are implemented along the spatial axis.
@@
where $s_{i}(t)$ is the state of site $i$ at time $t$ and $p$ is the bond probability of a connection between two adjacent sites. $z_{i}^{\pm}\in(0,1)$ is a random number taken from a uniform distribution. The system employs periodic boundary conditions along the spatial axis, as illustrated in Fig. 1. Each active site $s_{i}(t)$ propagates simultaneously to its left $s_{i-1}(t)$ and right $s_{i+1}(t)$ neighboring spatial positions during temporal evolution. The mutual independence of propagation probabilities $z_{i}^{\pm}\in(0,1)$ ensures the absence of spatial correlations. From the definitions above, the configuration of the system can be determined by the states of all sites.
The models of DP universality class also have their own independent critical exponents. The order parameter of the (1+1)-dimensional bond DP close to the critical point $p_{c}$ can be expressed by the steady-state particle density $\rho_{a}$ according to the critical exponent $\beta$ as [39],
@@
According to established literature, the theoretical critical threshold for (1+1)-dimensional bond percolation is $p_{c}=0.644700185(5)$ [54], determined through series analysis. Recent Monte Carlo simulations with high precision implementation yield $p_{c}=0.6447001(2)$ [55], obtained by single particle sources with $t=2^{14}$ time steps.
In the (1+1)-dimensional bond directed percolation (DP) model, there are two types of initial conditions for generating configurations. One is the homogeneous particle source, i.e., a fully occupied lattice where all spatial sites are initially occupied; the other is the single particle source, as only one site is initially occupied. Due to the reflection symmetry of the DP dynamics, both initial conditions lead to the same critical behavior, including the critical point and the critical exponents. In this work, we use the fully occupied lattice, as shown in Fig. 1 as input to the SNN algorithm. In the generated configurations, the first row represents the system at time $t=0$. Each row corresponds to spatial information at a given time step, and each column reflects the temporal evolution. To generate the configurations, we use periodic boundary conditions and originally start with fully occupied lattice (explained in our previous study [22]). The characteristic temporal length $t_{c}\sim L^{z/d}$ is set by the dynamical exponent $z=1.580(1)$ and the spatial dimension $d=1$, to ensure stable results [39].
2.2 The Ising model
(a) $T<T_{c}$
(b) $T=T_{c}$
(c) $T>T_{c}$
Figure 2: Configurations of the 2 dimensional Ising model corresponding to three different temperatures $T=1.52,2.269$ and $3.52$ respectively. The red lattice points represent spin-up states()($\uparrow$) and the blue lattice sites indicate spin-down states()($\downarrow$). The system configuration corresponds to a lattice size of L=32 with t=10000 time steps.
The Ising model [42, 43] serves as a paradigmatic model for systems exhibiting strong interactions between particles. Within the phenomenological framework, phase transitions can be characterized by order parameters, which assume non-zero values in ordered phases and vanish in disordered phases [56]. The two-dimensional Ising model considers a periodic lattice with exchange interactions limited to nearest neighbors. In the case of uniaxial magnetic materials, each lattice site can adopt either a spin-up or spin-down state. The Hamiltonian of the classical two-dimensional Ising model can be expressed as:
$H=-J\sum\limits_{<i,j>}\sigma_{i}\sigma_{j}-\sum\limits_{j}h_{j}\sigma_{j},$ (5)
where $\sigma_{i}\in\left\{-1,1\right\}$ represents the classical spin, $J$ denotes the interaction strength, and $h_{j}$ represents the external field. The critical temperature at which this model undergoes a phase transition is $T_{c}=2.269$ (for $h_{j}=0$)[42]. Fig. 2 depicts the configurations of the Ising model at three different temperatures. The selection of the ising model dataset for our study will be given in Section 3.2.
@@ 3.1 The Siamese Neural Network (SNN) method
Compared to other ML methods, the SNN employs a dual-branch architecture specifically designed to learn the dynamic similarity between pairs of configurations, making it particularly effective at capturing spatiotemporal correlations. It requires only “same-phase/different-phase” labels for configuration pairs—no explicit phase or critical-point labels for individual samples—resulting in much more efficient labeling, especially for ambiguous configurations near the critical region. Moreover, the same SNN architecture can be applied without modification to both equilibrium systems (e.g., the Ising model) and non-equilibrium systems (e.g., DP), demonstrating strong cross-system generalizability.
However, SNN also has its limitations. It outputs a similarity score rather than explicit class probabilities, which can make interpretation more challenging—an issue that has been noted in reviews of contrastive learning approaches. Its performance is highly sensitive to how representative the positive and negative sample pairs are; if these pairs are imbalanced or poorly chosen, training can suffer. Moreover, the deep and abstract nature of the learned feature space means that SNN embeddings often do not directly map to measurable physical quantities. Overall, while not a replacement for traditional or other ML methods, SNN offers a powerful complementary tool—especially for analyzing systems with complex spatiotemporal dynamics.
Figure 3: The structure of SNN. $S_{i}$ and $S_{j}$ represent a pair of input configurations (e.g., spatiotemporal configurations generated by bond probabilities $p_{i},p_{j}$ in the DP model, or spin configurations generated by temperatures $T_{i},T_{j}$ in the Ising model). Flatten denotes the operation of flattening a two-dimensional configuration into a one-dimensional vector. The sequence $\rightarrow\mathrm{BN}\rightarrow\mathrm{ReLU}$ indicates that the fully connected layer is followed by Batch Normalization and Rectified Linear Unit (ReLU) activation. The term $\|f_{1}(S_{i})-f_{2}(S_{j})\|$ denotes the feature difference between the two inputs. This difference is passed through Batch Normalization and ReLU activation, and then a Sigmoid activation function is applied to produce a similarity probability output.
@@
The specific SNN framework used in our work to study the phase transition of the (1 + 1) dimensional DP is shown in Fig. 3. For the better understanding, we name the first part of the SNN which produces $f(S_{i})$ as Feature extractor, and the later one estimates similarity as Similarity estimator. We use the raw configurations of $L=8,16,32,48$ and $64$ as input. The input is a set of data pairs $(S_{i},S_{j})$, $S_{i}$ enters one branch in the framework, and $S_{j}$ enters the other. The branches here are generated with the replica of the Feature extractor, sharing weight and the same hyper-parameters. $S_{i}$ and $S_{j}$ are two different configuration samples of the same type of data. Taking the branch with $S_{i}$ as an example, the input $S_{i}$ is flattened into one-dimensional data and fed into Feature extractor. $S_{i}$ flows firstly into a fully connected (FC) layer of $500$ neurons with a $ReLu$ activation function. Here a Batch Normalization (BN) is set to avoid over-fitting and speed up the NN to optimize its parameters before the $Relu$ activation function. One more FC layer in the Feature extractor maps the data flow into a feature representation $f_{1}$ of length $50$. $S_{i}\in\mathbb{R}^{L\times T}$ denote the input configuration pair, where $L$ is lattice size and $T$ is temporal dimension. The feature extraction process can be formalized as:
$f(S_{i})=W_{\text{feat}}\cdot h^{(1)}+b_{\text{feat}},$ (6)
with $h^{(1)}$ the hidden layer,
$h^{(1)}=\text{ReLU}\left(\text{BN}(W_{\text{flat}}\cdot\text{vec}(S_{i})+b_{\text{flat}})\right).$ (7)
Here $\text{vec}(\cdot)$ means a flatten operation, and $\text{BN}(\cdot)$ gives the Batch Normalization. $W_{\text{flat}}\in\mathbb{R}^{500\times(L\cdot T)}$ and $W_{\text{feat}}\in\mathbb{R}^{50\times 500}$ are the first and second FC layer weights, respectively. The entire calculation process of the first branch can be expressed as function $f_{1}(S_{i})$. Similarly, the second branch with $S_{j}$ has the same process, and the feature representation is recorded as $f_{2}(S_{j})$. Note that these two branches shares the weights for learning.
In the SNN, the weight update in the shared weight area is basically the same as in a single structure NN, except that the shared weights would update the same part in both branches simultaneously. Each branch of the SNN processes one input separately, and the forward propagation of the shared weight area calculates the same weights and biases. After that, the outputs of the two branches are merged to calculate a loss value. Once the loss is calculated, the gradient is determined through backpropagation. The gradient of the shared weight area is the sum of the gradients from the two branches. This shared weight mechanism ensures that the two branches remain consistent during training, thereby achieving effective feature extraction and comparison of similar or different inputs. Since the shared weights are the same in both branches, the weights are updated simultaneously.
@@
The shared weights scheme enable Feature extractor embeds data to the same space from any branches. Therefore, it allows to calculate the “distance" between two the feature representations $f_{1}(S_{i})$ and $f_{2}(S_{j})$. in order to compare the similarity of $S_{i}$ and $S_{j}$, here we define the “distance" by $L_{1}$ norm. Then the norm is fed into Similarity estimator. The $L_{1}$ norm emphasizes substantial differences in the learned representation space, analogous to the role of an order parameter in highlighting critical changes between phases. As a result, the model focuses on macroscopic statistical distinctions, rather than microscopic alignment. The network consists of two dense layer of FC with neurons $50$ and $1$ neurons and $Relu$ and sigmoidsigmoid activation function. Before the data flows get through the activation of the first layer, the flow is normalized with BN layer. The output of the Similarity estimator, $s(S_{i},S_{j})$, is converted into a probability of $[0,1]$ due to the $Sigmoid$ activation function, representing the similarity between $S_{i}$ and $S_{j}$. Then the similarity metric is computed as:
$s(S_{i},S_{j})=\sigma\bigg{(}W_{1\times 50}\cdot\text{ReLU}\bigg{(}\text{BN}(|f(S_{i})-f(S_{j})|)\bigg{)}+b_{50}\bigg{)},$ (8)
where $\sigma(\cdot)$ is sigmoid function.
@@ 3.2 Data sets of the SNN
For input data of the SNN, we use the full configurations of (1+1) dimension DP (time evolution step as the characteristic temporal length.). In each lattice size $L=8,16,32,48$ and $64$, we select $101$ values of $p\in[0,1]$ uniformly and generate $1000$ samples of configuration at each $p$. To minimize human intervention during training, we label the samples of configuration far away from the critical regime, i.e., the input configurations $S_{i}$ and $S_{j}$ are generated in the range of $p\in[0,0.1]\cup[0.9,1]$. The idea of labeling here is very different from the input of TL in Ref. [22]. The labels “1" or “0" indicates the input data pair $S_{i}$ and $S_{j}$ belong to the same or different phases to ensure maximum similarity or dissimilarity between them. Specifically, if $S_{i}$ and $S_{j}$ both belong to $[0,0.1]$ (or $[0.9,1]$), the input $(S_{i},S_{j})$ is labeled as “1". Or $S_{i}$ and $S_{j}$ are in the range of $[0,0.1]$ and $[0.9,1]$ respectively, the data pair is labeled as “0".
For the 2-dimensional Ising model, we employ the Wolff algorithm to generate the configurations. For each temperature point including those near the critical region($T\approx 2.269$) and in the low-temperature regime($T<T_{c}$), we perform 10000 Wolff cluster updates. The first $50\%$ of these steps (5000 steps) are exclusively used for thermalization, and sampling is conducted only during the remaining 5000 steps to eliminate the influence of the initial state. The temperature $T$ is selected from the range of 1.02 to 4.02, with an interval of 0.01, resulting in a total of 301 temperature points. For each temperature, 1000 samples are generated. The label set is chosen as $T\in[1.02,1.52]\cup[3.52,4.02]$. For training, the loss function is defined as:
@@
where $s$ is the similarity of ($S_{i}$, $S_{j}$) , and $y$ is the label, ranging from 0 to 1. The training epoch is set to $5000$, and our work in on python 3.10 and RTX 4090 Gpu platform with tensorflow 2.14.0 library.
For the process of test, a reference point, called "anchor" point, is selected to evaluate the similarity $s$ or dissimilarity (1s)(1-s) of all configurations. The output $s$ by the SNN represents the average similarity between configurations generated at specific bond probabilities (for DP) or temperatures (for Ising) and anchor $p_{a}$. For a given pair $(S_{i},S_{j})$, the SNN computes their similarity score $s(S_{i},S_{j})\in[0,1]$. When analyzing similarity across parameter ranges, the reported curve is the average over all pairwise comparisons at each parameter value. This approach allows the SNN to capture the statistical similarity between configurations, even when individual samples may appear random due to fluctuations near criticality. By averaging over multiple configurations, the SNN effectively smooths out noise and highlights the underlying phase-dependent patterns. The anchor $p_{a}$ can be selected arbitrarily within the range of $p\in[0,1]$. By taking the samples of configuration at the anchor $p_{a}$ as $S_{j}$ and the samples of the training set as $S_{i}$, in this way the similarity between the samples at the anchor and the training samples can be estimated through the SNN. And in detail the dependence of SNN results on anchor point will be discussed in Section 4.2. Then the output of SNN, $P_{0}$, can be expressed as two curves within bond probability $p\in[0,1]$: the curve $s$ of similarity and 1s1-s of dissimilarity. The intersection of these two curves or $p$ at $P_{0}=50\%$ represents the critical point of phase transition in the system.